# Networks with many structural scales: a Renormalization Group perspective

### v2

@@ Abstract
Figure 2: Clustered networks. RW heat capacity versus diffusion time for: (a) $(1,2)$ flowers with $s=6,8,9,10$ and $11$, hierarchical levels (b) $(2,2)$ flowers with $s=5,6,7$, and $8$. For $(u,v)$ flowers, the number of nodes is $N_{s}=\left(\frac{w}{w-1}+\frac{w-2}{w-1}w^{s}\right)$, with $w=u+v$. (c) KH with $m=2$ and (d) KH with $m=3$ of sizes $N=\{1,2.5,4,8\}\times 10^{3}$. Insets show the scaling $\tau_{F}$ as a function of the system size. Orange dashed lines and solid black lines represent the theoretical expectation. All curves have been averaged over $10^{3}$ independent realizations.
@@
Fig.2(b) sheds light on their differences: $(u,v)$ flowers with $u>1$ exhibit characteristic oscillations emerging from the discrete nature of the recursive network growth process (much as log-periodic oscillations in standard fractal analyses of deterministic structures, as stem from their discrete-scale invariance Sornette (2006); Vicsek et al. (1994)). Thus, we demonstrate that the method detects the presence of discrete scale invariance on networks.
@@
Third, to explore whether self-similarity can exist in stochastic scale-free graphs with small-world properties Rozenfeld et al. (2007), we consider the Kim and Holme (KH) model Holme and Kim (2002). This requires modifying BA networks by introducing a probability $p$ for each newly added node to form a triangle with existing nodes (see SM SM ). The resulting networks show a power-law degree distribution with high clustering coefficient. LRG analysis of KH networks reveals scale-invariance only when $p=1$. As shown in Figs.2(c) and (d), $C_{RW}$ (and $C_{L}$ SM ) shows a plateau with spectral dimension from $d_{s}=2.57(1)$ to $d_{s}=3.65(1)$ as $m$ increases (see SM SM ).
@@
Fourth, we have investigated the behavior of networks with a built-in hierarchical structure. In particular, we have analyzed the Dyson graph Dyson (1969); Agliari and Tavani (2017): a fully-connected deterministic graph with hierarchically organized link weights, resulting in a tunable spectral dimension $d_{s}=2/(2\sigma-1)$ where $\sigma$ is a scaling parameter controlling the weight strength at every scale (see SM SM ). Fig.3(a) shows the constant plateau of $C_{L}$ for a Dyson network with $\sigma=0.85$, consistent with the theoretical value. We have also examined hierarchical modular networks (HMNs) that were proposed using inspiration from actual brain networks Moretti and Muñoz (2013). In HMNs, nodes grouped in basal fully-connected modules are recursively coupled with nodes in other moduli, establishing inter-modular links randomly in a hierarchical fractal-like manner (see SM SM ). Fig.3(b) shows their heat capacity: HMNs are scale-invariant for any set of parameters, with spectral dimension in the range $d_{s}\in(1.25,2)$ (see SM SM ).
@@
Figure 3: Hierarchical networks. Heat capacity versus diffusion time for: (a) Dyson graphs with $\sigma=0.85$ and $s=11,13,14,15$ hierarchical levels. The network size is $N=2^{s}$ in both cases. The upper left inset shows the weighted adjacency matrix. (b) HMNs with $m_{0}=3$ and $\alpha=2$ (see SM SM ) for $s=9,10,11,12$ hierarchical levels ($N=2^{s}m_{0}$). Insets show the scaling of $\tau_{F}$ as a function of $N$. Orange dashed and solid black lines represent the theoretical expectation. (c) The HC for different thresholds (see legend). (d) The scaling of $\tau_{F}$ for coarse-grained versions of the HC with $T=3\cdot 10^{-2}$ and different fixed values of $\tau=10^{-1}$, $10^{0}$, and $10^{1}$. Black dashed line corresponds to $d_{s}=1.9$.
@@
Finally, we have analyzed the Human Connectome (HC) structural brain network Fornito et al. (2016) (see SM SM ). The analysis of $C_{L}$ for the largest available (thresholded) HC network is shown in Fig.3(c), while Fig.3(d) shows the finite-size scaling of $\lambda_{F}$ for reduced versions of the HC. Both analyses support an underlying scale-invariant topology with spectral dimension $d_{s}\sim 1.9(1)$.
@@
The observation of scale-invariance across length and time scales led to the development of RG ideas Kardar (2007); Amit and Martin-Mayor (2005) whose initial implementation took advantage of the geometrical homogeneity and translational invariance to perform scaling of statistical physical systems, computing critical exponents Wilson and Fisher (1972), and leading to a categorization of apparently diverse phase transitions into a few universality classes Hohenberg and Halperin (1977); Ódor (2004).
Extending universality to networks with heterogeneous and non-local connectivity properties has long challenged physicists. Despite remarkable advances have been recently made Song et al. (2005); Serrano et al. (2008); Rozenfeld et al. (2010); García-Pérez et al. (2018); Zheng et al. (2020), fully understanding scale-invariance in generic networks remains unresolved. Initially, network scale-invariance was associated with power-law degree distributions Serafino et al. (2021). However, as we explicitly show, scale-freeness of node degrees is neither necessary nor sufficient for true self-similarity Broido and Clauset (2019). Instead, the LRG Villegas et al. (2023a, 2022), provides a general framework to characterize and classify non-trivial structural LRG fixed points, describing their scaling properties and offering a classification in terms of universality classes. A constant heat-capacity indicates the presence of scale-invariance or self-similarity in graphs, characterized by a constant entropy-loss rate during network coarse-graining. We have linked the constant entropy-loss rate to the network spectral dimension, $d_{s}$ —computed using either $\hat{L}$ or $\hat{L}_{RW}$— confirming its role as a natural generalization of the Euclidean dimension Cassi (1992); Cassi and Fabbian (1999); Burioni and Cassi (1996).
Our LRG analysis identifies regular lattices, trees, clustered networks, and hierarchical networks as fundamental classes of scale-invariant networks. Beyond regular lattices, we highlight (u-v)-flowers and KH networks as unique structures combining local clustering with hubs at all scales to generate self-similar topologies. Trees are bound to $d_{s}\leq 2$, and simply adding hubs appears insufficient to achieve higher-dimension values. Generally, as the dimension increases, it becomes more challenging to generate self-similar heterogeneous networks Rozenfeld et al. (2007). Our findings are promising for studying biological and socio-technological networks; notably, hierarchical networks, including the HC Sporns et al. (2004); Hagmann et al. (2008) which exhibit robust scale invariance with $d_{s}\approx 1.9(1)$, corroborating previous results Zheng et al. (2020). Our findings provide a solid ground for future dynamical RG theoretical calculations on these structures and for identifying new signatures of scale invariance in real-world networks.
@@ Acknowledgments
The authors thank T.Gili, D. Cassi, and R. Burioni for very useful discussions and comments. P.V. and M.A.M. acknowledge financial support from the Spanish "Ministerio de Ciencia, Innovación y Universidades" and the "Agencia Estatal de Investigación (AEI)" under Project Ref. PID2020-113681GB-I00 funded by MICIN/AEI/10.13039/501100011033.

### v3

@@ Abstract
Figure 2: Clustered networks. $C_{RW}$ as a function of the resolution scale $\tau$ for: (a) $(1,2)$ flowers with $s=6,8,9,10$ and $11$, hierarchical levels, respectively. (b) $(2,2)$ flowers with $s=5,6,7$, and $8$ levels resp. For $(u,v)$ flowers, the number of nodes is $N_{s}=\left(\frac{w}{w-1}+\frac{w-2}{w-1}w^{s}\right)$, with $w=u+v$. (c) KH networks with $m=2$ and (d) KH with $m=3$ of sizes $N=\{1,2.5,4,8\}\times 10^{3}$. Insets show the scaling of $\tau_{F}$ as a function of network size $N$. Orange dashed lines and solid black lines represent the theoretical expectation. All curves have been averaged over $10^{3}$ independent realizations.
@@
Fig.2(b) illustrates that $(u,v)$ flowers with $u>1$ exhibit characteristic oscillations emerging from the discrete nature of the recursive network growth process (as log-periodic oscillations stemming from discrete-scale invariance in standard fractal analyses Sornette (2006); Vicsek et al. (1994)). This demonstrates that the RG method is also able to detect the presence of discrete scale invariance on networks.
@@
Third, to explore self-similarity in stochastic scale-free graphs with small-world properties Rozenfeld et al. (2007), we consider the Kim-Holme (KH) network model Holme and Kim (2002). This modifies BA networks by introducing a probability $p$ for each newly added node to form a triangle with existing nodes (see SM SM ). The resulting networks show a power-law degree distribution and a high clustering coefficient. LRG analysis of KH networks reveals scale-invariance only for $p=1$. As shown in Figs.2(c) and (d), $C_{RW}$ (and $C_{L}$ SM ) shows a plateau with spectral dimension growing with $m$ from $d_{s}=2.57(1)$ to $d_{s}=3.65(1)$ (see SM SM ).
@@
Fourth, we investigate networks with a built-in hierarchical structure. In particular, we analyze the Dyson graph Dyson (1969); Agliari and Tavani (2017): a fully-connected deterministic graph with hierarchically organized link weights, resulting in a tunable spectral dimension $d_{s}=2/(2\sigma-1)$, being $\sigma$ a scaling parameter controlling the weight strength at every scale (see SM SM ). Fig.3(a) shows the constant plateau of $C_{L}$ for a Dyson graph with $\sigma=0.85$, consistent with the theoretical value. We have also examined hierarchical modular networks (HMNs) that were proposed using inspiration from brain networks Moretti and Muñoz (2013). In HMNs, nodes grouped in basal fully-connected modules are recursively coupled with nodes in other moduli, establishing inter-modular links randomly in a hierarchical fractal-like manner (see SM SM ). Fig.3(b) shows their heat capacity: HMNs are scale-invariant for any set of parameters, with spectral dimension in the range $d_{s}\in(1.25,2)$ (see SM SM ).
@@
Figure 3: Hierarchical networks. Heat capacity as a function of the resolution scale $\tau$ for: (a) Dyson graphs with $\sigma=0.85$ and $s=11,13,14,15$ hierarchical levels resp.(network size is $N=2^{s}$ and the upper left inset shows the weighted adjacency matrix); (b) HMNs with $m_{0}=3$ and $\alpha=2$ (see SM SM ) for $s=9,10,11,12$ hierarchical levels resp. ($N=2^{s}m_{0}$). Right insets (a) and (b) show the scaling of $\tau_{F}$ as a function of $N$. Orange dashed and solid black lines, respectively, show the theoretical expectation. (c) HC network for different thresholds (see legend). (d) Scaling of $\lambda_{F}$ for coarse-grained versions of the HC with $T=3\cdot 10^{-2}$ and $\tau-$values of $10^{-1}$, $10^{0}$, and $10^{1}$, which are superimposed. Black dashed line represents the scaling using $d_{s}=1.9$.
@@
Finally, we have analyzed the Human Connectome (HC) structural brain network Fornito et al. (2016) (see SM SM ). To account for fluctuations in single realizations of empirical networks, we apply the LRG Villegas et al. (2023a) to analyze reduced HC versions, performing finite-size scaling (see SM for the detailed procedure SM ). Due to the high density of weak links, the HC must be sparsified by removing links below a threshold (T) to reveal non-trivial scale-invariant features, but avoiding the trivial RT expected at the percolation critical point (at $T\approx 0.06$ in this case). Rusults of the G analysis of $C_{L}$ for thresholded HC networks is shown in Fig.3(c), while Fig.3(d) illustrates the robust finite-size scaling of $\lambda_{F}$ for reduced versions of a HC at different $\tau$-values. Both analyses support an underlying scale-invariant topology with spectral dimension $d_{s}\sim 1.9(1)$ (in agreement with Zheng et al. (2020)).
@@
The observation of scale-invariance across length and time scales led to the development of RG ideas Kardar (2007); Amit and Martin-Mayor (2005) whose initial implementation took advantage of the geometrical homogeneity of interactions and translational invariance of the embedding space to perform scaling of statistical physical systems, computing critical exponents Wilson and Fisher (1972), and leading to classify apparently diverse phase transitions into a few universality classes Hohenberg and Halperin (1977); Ódor (2004).
Extending universality to networks with heterogeneous and non-local connectivity properties has long challenged physicists. Despite remarkable advances have been recently made Song et al. (2005); Serrano et al. (2008); Rozenfeld et al. (2010); García-Pérez et al. (2018); Zheng et al. (2020), fully understanding scale-invariance in generic networks remains unresolved. Initially, network scale-invariance was associated with power-law degree distributions Serafino et al. (2021). However, as we explicitly show, scale-freeness of node degrees is neither necessary nor sufficient for true self-similarity Broido and Clauset (2019).
Instead, the LRG Villegas et al. (2023a, 2022) provides a framework capable of characterizing and classifying non-trivial structural fixed points, describing their scaling properties and offering a classification in terms of universality classes. A constant heat-capacity indicates the presence of scale-invariance or self-similarity in graphs, characterized by a constant entropy-loss rate during network coarse-graining. We have linked the constant entropy-loss rate to the network spectral dimension, $d_{s}$ —computed using either $\hat{L}$ or $\hat{L}_{RW}$— confirming its role as a natural generalization of the Euclidean dimension Cassi (1992); Cassi and Fabbian (1999); Burioni and Cassi (1996).
Our LRG analysis identifies regular lattices, trees, clustered networks, and hierarchical networks as fundamental classes of scale-invariant networks. Beyond regular lattices, we highlight (u-v)-flowers and KH networks as unique structures combining local clustering with hubs at all scales to generate self-similar topologies. Trees are bound to $d_{s}\leq 2$, and simply adding hubs appears insufficient to achieve large dimension values. Generally, as the dimension increases, it becomes more challenging to generate self-similar heterogeneous networks Rozenfeld et al. (2007). Our findings are promising for studying biological and socio-technological networks; notably, hierarchical networks, as the HC Sporns et al. (2004); Hagmann et al. (2008) which exhibit robust scale invariance with $d_{s}\approx 1.9(1)$, corroborating previous results Zheng et al. (2020). We provide a solid ground for future dynamical RG theoretical calculations on these structures and for identifying new signatures of scale invariance in real networks.
@@ Acknowledgments
We thank T. Gili, D. Cassi, and R. Burioni for very useful discussions and comments. P.V. and M.A.M. acknowledge financial support from the Spanish "Ministerio de Ciencia, Innovación y Universidades" and the "Agencia Estatal de Investigación (AEI)" under Project Ref. PID2020-113681GB-I00 funded by MICIN/AEI/10.13039/501100011033 and Grant No. PID2023-149174NB-I00 financed by MICIU/AEI/10.13039/501100011033 and EDRF/EU funds.